# Wireless 6G Connectivity for Massive Number of Devices and Critical Services

### v2

@@ I Introduction
One of the defining features of 5G was the native support for Internet of Things (IoT) through massive machine-type communication (mMTC) and ultra-reliable low-latency communications (URLLC). While mMTC aimed to provide delay-tolerant service for a very large number of low-cost devices, URLLC targeted human-centric and machine-centric real-time applications, such as virtual reality and self-driving cars. Along with enhanced mobile broadband (eMBB), these services constituted the “5G triangle,” which served as the guiding framework for the 5G research and standardization efforts [1]. At the conception of 5G, it had been thought that mMTC and URLLC, as well as their combinations with eMBB were sufficient to cover almost all IoT use cases. Nevertheless, with the evolution and deployment of 5G it started to emerge that the requirements for mMTC and URLLC would need to be redefined, as illustrated in Fig. 1. For instance, the stringent latency requirements of URLLC can be relaxed to a set of more general timing requirements [2]. Furthermore, new use cases, especially centered around machine learning (ML) and artificial intelligence (AI), have introduced new scenarios that were unthinkable when the 5G vision was initially sketched. Finally, there is a major change in the overall access infrastructure by embracing Non-Terrestrial Networks (NTN) [3, 4]. These networks introduce moving elements in the infrastructure, such as LEO satellites and significantly augment the spatial coverage, thereby the number of potentially accessing devices. Although NTNs will have an important role in massive connectivity, as also indicated in Fig. 0(b), in this article we will not cover the NTN aspects in detail, as it would require a more elaborate discussion on the specifics of the related network architecture, which is significantly different from the terrestrial one.
@@
As we move towards 6G, the 5G triangle has been expanded to a hexagon, defined by enhanced variants of the three 5G service types, as well as new scenarios targeting integrated sensing and communication, integrated AI and communication, and ubiquitous connectivity [5]. With the four overarching requirements of sustainability, security/privacy/resilience, ubiquitous intelligence, and connecting the unconnected, the support for IoT is expected to play an even larger role in 6G. To fulfill the long-term 6G vision that involves IoT connectivity requires, however, progress in three key areas: communication for massive number of devices, communication for critical services, and service coexistence and slicing capabilities.
@@
The evolution of massive communication from 5G to 6G will be characterized primarily by two aspects. The first aspect is the even higher device densities targeted in 6G, ranging from $10^{6}$ to $10^{8}$ devices/km2 [5]. This requires more scalable random access protocols that can resolve transmissions from a large number of uncoordinated, simultaneously transmitting devices. The second aspect is a widening of the device spectrum, ranging from zero-energy devices with extremely limited capabilities and long lifetimes, to highly capable devices equipped with advanced sensors, such as microphones and cameras, and able to execute ML algorithms locally or in the cloud. ML-driven decisions will increasingly govern the traffic patterns and determine the communication targets, since the need to communicate will be triggered by unpredicted events rather than occurring periodically. In addition, advanced devices require frequent firmware updates, and may continuously participate in distributed learning algorithms, such as federated learning, leading to closed-loop communication. In conclusion, massive connectivity in 6G will shift the focus from uplink-oriented massive IoT to a wide range of traffic patterns including both small, sporadic transmissions, and closed-loop uplink and downlink driven by interactive applications and distributed learning.
@@
Critical communication refers to communication with strict performance requirements, and was introduced in 5G as URLLC. The aim was to provide a service that could replace traditional cabled industrial networks by delivering packets within milliseconds with a reliability in the order of $1-10^{-5}$ (five nines) or higher. While the vision in 6G is to provide even stricter latency and reliability requirements [5], meeting the vision of critical communication requires fundamental changes. While this paper is focused only on the wireless access part, in a broader perspective, timing and reliability need to be considered end-to-end at the system level, including the mobile core network and various hardware components that influence the overall reliability. Furthermore, since the initial vision of 5G was defined, there has been a realization that the latency and reliability can be decoupled to some extent by focusing on the specific task and goals of the application rather than the link-level requirements [2]. However, maximizing these benefits requires a close integration between the communication system and the application itself, and identifying the control interface will be an important task in the development of 6G. Meeting the complex demands of communication systems, closely linked with control systems and stringent constraints on delay, reliability, availability and resilience, necessitates the development of advanced AI techniques, incorporating optimization theory, explainable AI, generative AI and digital twin, for optimal performance.
@@ II-A Massive connectivity in 6G: A mix of uplink and downlink
The fundamental problem in massive connectivity is to serve a large number, $N$, of users that access the channel in a sporadic fashion, as illustrated in Fig. 2. Specifically, only a small subset, comprising $K\ll N$ users, are active at any given time, and the set of active users is unknown a priori. Considering the target device density in 6G, $K$ is often assumed to be at most a few hundred, whereas $N$, depending on the scenario, can be in the order of thousands or even millions.
@@
In 5G, massive connectivity was introduced as the mMTC use case and enabled by the NB-IoT and LTE-M technologies, which provided native support for low-power devices and short packet transmissions through simplified connection establishment procedures [7]. However, while 5G mMTC was primarily targeting uplink traffic [8], new use cases envisioned in 6G, such as cyber-physical systems, interactive IoT, and distributed inference/learning, will bring more symmetric traffic patterns that require equal uplink and downlink, leading to massive closed-loop communication [9, 10].
@@
In the following, we review the current directions and techniques for achieving massive connectivity in 6G. We first present an overview of grant-based, grant-free, and unsourced random access as the main categories of massive uplink. We then go into details with sparse recovery as a promising building block for efficient and scalable uplink connectivity. Finally, we discuss recent work towards designing massive connectivity for the downlink.
@@
II-B Grant-Based, Grant-Free, and Unsourced Random Access for Massive Uplink Connectivity
The sporadic activity in the uplink leads to the random access problem, in which the set of transmitting users is unknown a priori to both the transmitters and the base station. Although uplink random access is part of most commercially available mMTC technologies, such as those in 5G, they are implemented using inefficient techniques that are likely to fall short of the scale envisioned in 6G massive connectivity. As a result, scaling random access to a massive number of users is an ongoing research task, considered in both academia and industry, which has led to a number of recent developments. Broadly speaking, these developments can be divided into grant-based, grant-free, and unsourced random access.
@@ II-B Grant-Based, Grant-Free, and Unsourced Random Access for Massive Uplink Connectivity
Figure 3: Grant-based and grant-free random access. In grant-based random access (2(a)), the preamble is transmitted first, allowing the base station to assign dedicated resources for the data transmission (i.e., a grant). In grant-free random access (2(b)), the preamble and data are transmitted without intermediate feedback from the base station, typically in a sequential fashion.
@@ II-B1 Grant-Based Random Access
Both NB-IoT and LTE-M use orthogonal preambles in the first step, where each active user transmits a preamble randomly selected from a pool of orthogonal preambles. While this leads to low complexity at the decoder, the coherence time of the channel limits the number of available preambles, which leads to a high collision probability and subsequent packet loss. Consequently, it is well known that the throughput measured by the fraction of non-colliding transmissions is upper bounded by $1/e\approx 0.37$. Combined with the low number of preambles, this severely limits the number of users that can be supported by the system.
The performance of grant-based schemes can be greatly improved by using more sophisticated preamble decoding techniques, such as sparse recovery algorithms that rely on compressed sensing [11, 12]. By exploiting the sparsity of the transmitted preambles, the orthogonality requirement can be relaxed, which significantly increases the number of available preambles without increasing the complexity of the transmitter. The number of available preambles needs to be selected to balance the collision probability, the probability of false alarms, and the complexity of the activity detection algorithm. When the total number of users is not too large, the error is dominated by collisions rather than false alarms, and it may be advantageous to assign a unique preamble to each user to completely avoid collisions. Furthermore, this enables user identification “for free”, since the transmitting users can be identified based from the decoded preambles, and it is even possible to embed a few bits of information in the preambles as well, e.g., to indicate the amount of data that the user has to transmit [13]. This approach can be further justified by the result that the probability of activity detection error can theoretically be made to go to zero as the number of base station antennas goes to infinity [12]. However, when the total number of users is large, the computational complexity of the activity detection algorithm may be prohibitive, and false alarms may also be more prominent than collisions. In this case, letting the users randomly select their preamble may lead to better overall performance-complexity tradeoff. We return to the use of sparse recovery for massive uplink in Sec. II-C.
@@ II-B2 Grant-Free Random Access
The main disadvantage of grant-based random access is the latency and power consumption introduced by the three protocol steps. Furthermore, when the packets are short, the additional overhead introduced by the grant transmission may outweigh the advantage of having dedicated channel resources for the packet transmissions. An alternative strategy that avoids this procedure is grant-free random access, illustrated in Fig. 2(b), where the active devices transmit their preambles and packets immediately, typically in a sequential manner, without waiting for intermediate feedback from the base station. The recently-added Small Data Transmission feature in 5G is an example of steps in this direction. To decode the transmitted packets, the base station performs activity detection and channel estimation based on the preambles, and then decodes the messages coherently while treating the interfering transmissions as noise.
@@ II-B3 Unsourced Random Access
The majority of the protocols developed for unsourced random access rely on sparse recovery, possibly combined with additional coding [18, 19], although techniques based on, e.g., tensor decomposition have been proposed as well [20]. Furthermore, many random access schemes designed for grant-free access can be made unsourced by making the preamble selection random or determined based on the message data (as opposed to assigning a unique preamble to each user).
@@ II-C Sparse Recovery for Massive Random Access
Regardless of the type of random access, escaping the high collision probability associated with the use of orthogonal preamble transmissions is necessary to substantially increase the number of users that can be supported by the random access procedures. As a central component towards realizing this goal, this section provides a brief review of the foundations behind sparse recovery for massive random access. To simplify the presentation, we will focus on the case where each user is assigned a unique preamble, which precludes collisions. We consider a scenario with $N$ single-antenna devices and a single base station with $M$ antennas. Let $\mathbf{s}_{k}\in\mathbb{C}^{L\times 1}$ denote the preamble assigned to user $k$, $k=1,\ldots,N$, which is assumed to be known at the base station. Suppose a random number $K$ out of the $N$ users are active, and let $a_{k}=1$ if user $k$ is active, otherwise $a_{k}=0$. The channel between device $k$ and the base station is assumed to be given by the random vector $g_{k}\mathbf{h}_{k}\in\mathbb{C}^{M\times 1}$, composed by large scale fading coefficient $g_{k}\in\mathbb{R}$, and unknown Rayleigh fading coefficients $\mathbf{h}_{k}\in\mathbb{C}^{M}$ drawn independently from $\mathcal{CN}(0,1)$. The signal $\mathbf{Y}\in\mathbb{C}^{L\times M}$ received by the base station can then be written as

### v3

@@ I Introduction
One of the defining features of 5G was the native support for Internet of Things (IoT) through massive machine-type communication (mMTC) and ultra-reliable low-latency communications (URLLC). While mMTC aimed to provide delay-tolerant service for a very large number of low-cost devices, URLLC targeted human-centric and machine-centric real-time applications, such as virtual reality and self-driving cars. Along with enhanced mobile broadband (eMBB), these services constituted the “5G triangle,” which served as the guiding framework for the 5G research and standardization efforts [1]. At the conception of 5G, it had been thought that mMTC and URLLC, as well as their combinations with eMBB were sufficient to cover almost all IoT use cases. Nevertheless, with the evolution and deployment of 5G it started to emerge that the requirements for mMTC and URLLC would need to be redefined, as illustrated in Fig. 1. For instance, the stringent latency requirements of URLLC can be relaxed to a set of more general timing requirements that dynamically adjust to the current state of the system, such as Age of Information (AoI), Age of Loop (AoL), and Value of Information (VoI) [2]. Furthermore, new use cases, especially centered around machine learning (ML) and artificial intelligence (AI), have introduced new scenarios that were unthinkable when the 5G vision was initially sketched, such as federated learning and intelligent digital twins. Finally, there is a major change in the overall access infrastructure by embracing Non-Terrestrial Networks (NTN) [3, 4]. These networks introduce moving elements in the infrastructure, such as LEO satellites and significantly augment the spatial coverage, thereby the number of potentially accessing devices. Although NTNs will have an important role in massive connectivity, as also indicated in Fig. 1(b), in this article we will not cover the NTN aspects in detail, as it would require a more elaborate discussion on the specifics of the related network architecture, which is significantly different from the terrestrial one.
@@
As we move towards 6G, the 5G triangle has been expanded to a hexagon, defined by enhanced variants of the three 5G service types, as well as new scenarios targeting integrated sensing and communication, integrated AI and communication, and ubiquitous connectivity [5]. With the four overarching requirements of sustainability, security/privacy/resilience, ubiquitous intelligence, and connecting the unconnected, the support for IoT is expected to play an even larger role in 6G. To fulfill the long-term 6G vision that involves IoT connectivity requires, however, progress in three key areas: communication for massive number of devices, communication for critical services, and service coexistence and slicing capabilities. It is important to outline that the denomination “6G” can be understood in two different ways. In a broader sense, 6G refers to the future wireless connectivity, accompanied with sensing, positioning, learning, etc. In a narrower sense, 6G refers to the next standardized generation of wireless cellular communication system. The discussion on 6G in this paper should be understood in the broader sense, but in specific parts, such as Section II-E, we elaborate upon the potential impact on standardization.
@@
The evolution of massive communication from 5G to 6G will be characterized primarily by two aspects. The first aspect is the even higher device densities targeted in 6G, ranging from $10^{6}$ to $10^{8}$ devices/km2 [5]. This requires more scalable random access protocols that can resolve transmissions from a large number of uncoordinated, simultaneously transmitting devices. The second aspect is a widening of the device spectrum, ranging from zero-energy devices with extremely limited capabilities and long lifetimes, to highly capable devices equipped with advanced sensors, such as microphones and cameras, and able to execute ML algorithms locally or in the cloud. By predicting when and what data are relevant at a given time, ML-driven decisions will increasingly govern the traffic patterns and determine the communication targets, since the need to communicate will be triggered by unpredicted events rather than occurring periodically. For example, an AI may autonomously request data from IoT devices based on what is needed to reduce the uncertainty of a digital twin model, or because it predicts that the data will soon be requested by an external entity. Besides data requests, advanced devices also require frequent firmware updates, and may continuously participate in distributed learning algorithms, such as federated learning, leading to closed-loop communication. In conclusion, massive connectivity in 6G will shift the focus from uplink-oriented massive IoT to a wide range of traffic patterns including both small, sporadic transmissions, and closed-loop uplink and downlink driven by interactive applications and distributed learning.
@@
Critical communication refers to communication with strict performance requirements, and was introduced in 5G as URLLC. The aim was to provide a service that could replace traditional cabled industrial networks by delivering packets within milliseconds with a reliability in the order of $1-10^{-5}$ (five nines) or higher. While the vision in 6G is to provide even stricter latency and reliability requirements [5], meeting the vision of critical communication requires fundamental changes. While this paper is focused only on the wireless access part, in a broader perspective, timing and reliability need to be considered end-to-end at the system level, including the mobile core network and various hardware components that influence the overall reliability. Furthermore, since the initial vision of 5G was defined, there has been a realization that the latency and reliability can be decoupled to some extent by focusing on the specific task and goals of the application rather than the link-level requirements [2]. However, maximizing these benefits requires a close integration between the communication system and the application itself, and identifying the control interface will be an important task in the development of 6G. Meeting the complex demands of communication systems, closely linked with control systems and stringent constraints on delay, reliability, availability and resilience, necessitates the development of advanced AI techniques, incorporating optimization theory, explainable AI, generative AI and digital twin, for optimal performance. For instance, AI can help bridging the gap between model accuracy and complexity in interacting systems comprising both wireless channels and physical sensors and actuators, and where the relation between inputs and outputs has no simple mathematical expression. This is particularly important when the tail behavior of the model predictions are crucial, as in systems that require high reliability.
@@ II-A Massive Connectivity in 6G: A Mix of Uplink and Downlink
Massive connectivity refers to the problem of serving a very large number of users that access the channel in a sporadic fashion, as illustrated in Fig. 2. Traditionally, massive connectivity has been associated with simple IoT devices, such as sensors, that wake up periodically or sporadically to transmit short sensor readings to the base station. This was also the motivation in 5G, where massive connectivity was introduced as the mMTC use case and enabled by the NB-IoT and LTE-M technologies, which provided native support for low-power devices and short packet transmissions through simplified connection establishment procedures [7].
@@
Massive connectivity will remain a central component of 6G, where the vision is to support even higher device densities and serving a broader spectrum of device capabilities. In particular, data collection and environmental monitoring will be crucial for important 6G applications, such as AI-based predictions and decision making, which may involve devices that make coordinated, intelligent decisions, such as actively requesting information or initiating model retraining. Some important applications of massive connectivity in 6G include:
  • Sensing and actuation for digital twins: Digital twinning is being promoted as one of the central use cases in 6G, where physical objects will have virtual representations that can be used for monitoring, simulations, or virtual interactions in, e.g., industrial and smart traffic scenarios [8]. Massive connectivity plays a key role in enabling the data collection that will allow the digital twins to accurately represent their physical counterparts in real-time. The collected data may then be fused with data from other sensors or combined with, e.g., sensing and localization information collected by the base station within the same temporal window [9]. Finally, the simulations and interactions with the digital twins may eventually trigger actions or decisions that need to be communicated back to the devices for execution.
  • Logistics: The ubiquitous connectivity of 6G will enable IoT-based worldwide tracking of objects for logistics and supply chain management [10]. Depending on the complexity of the IoT device, applications may span from simple location reporting to tamper-resistant logging using distributed ledger technology, such as blockchain [5].
  • Massive sensing using zero-energy devices: Ultra simple, low-cost, battery-less devices that harvest energy from, e.g., vibrations or radio waves drastically simplifies the deployment of IoT. Without the need of a battery, these devices can be made extremely small and printed on things like textiles and warehouse packages for tracking and sensing purposes [10, 11].
  • Channel access in densely populated areas: 6G aims to target even higher device densities compared to 5G, such as crowded stadiums [5]. In such scenarios, allocating dedicated resources to each device for coordination of, e.g., scheduling requests, is impossible, and efficient and scalable massive connectivity technologies are crucial to ensure a good user experience.
While the overhead reductions introduced in 5G mMTC were crucial to enable the support of IoT, the access procedures in 5G still perform significantly worse in terms of energy and spectral efficiency compared to what is possible [12]. Consequently, 6G will need additional technological innovations to support the envisioned target device densities and applications in 6G without imposing a significant cost in terms of spectral efficiency. In this respect, one central protocol component that can benefit from substantial improvements is the random access procedure, which is needed due to the sporadic uplink transmissions, since the set of transmitting users is unknown a priori to both the transmitters and the base station. Most commercially available mMTC technologies, including those in 5G, are still implemented using techniques that resemble the first random access protocol, namely slotted ALOHA [13]. While slotted ALOHA protocols are simple to implement, it is well known that they have significant limitations in terms of energy and spectral efficiency, which has also been documented in the context of 5G [12]. Furthermore, since random access is central in any massive connectivity protocol regardless of its specific requirements, its performance will be reflected across all applications. Besides random access, several of the new applications envisioned in 6G, such as interactive digital twins, blockchain-enabled logistics, and inference/learning, will bring more symmetric traffic patterns that require equal uplink and downlink, leading to massive closed-loop communication [10, 14]. Since 5G mMTC was primarily targeting uplink traffic [15], the downlink channel for IoT is another area that will require new innovations in 6G.
@@
In the remainder of this section, we review the enabling technologies for enhancing the massive connectivity support in 6G. We first present an overview of the technologies in the context of grant-based, grant-free, and unsourced random access as the main categories of massive uplink. We then go into details with sparse recovery as a promising and general building block for efficient and scalable uplink connectivity. This is followed by an overview on recent work towards designing massive connectivity for the downlink. We conclude the section by discussing what is needed to integrate these developments into 6G standards and outlining some remaining open problems.
@@
II-B Grant-Based, Grant-Free, and Unsourced Random Access for Massive Uplink Connectivity for 6G
Broadly speaking, the recent developments in massive random access protocols can be divided into grant-based, grant-free, and unsourced random access. We discuss each category in turn.
@@ II-B Grant-Based, Grant-Free, and Unsourced Random Access for Massive Uplink Connectivity for 6G
Figure 3: Grant-based and grant-free random access. In grant-based random access (3(a)), the preamble is transmitted first, allowing the base station to assign dedicated resources for the data transmission (i.e., a grant). In grant-free random access (3(b)), the preamble and data are transmitted without intermediate feedback from the base station, typically in a sequential fashion.
@@ II-B1 Grant-Based Random Access
Both NB-IoT and LTE-M use orthogonal preambles in the first step, where each active user transmits a preamble randomly selected from a pool of orthogonal preambles without the possibility of multipacket reception. While this leads to low complexity at the decoder, it is well known that the throughput measured by the fraction of non-colliding transmissions is upper bounded by $1/e\approx 0.37$ [13]. Consequently, the utilization of the channel resources (i.e., the spectral efficiency) is inherently low. Furthermore, since the number of available preambles is limited by the channel coherence length, the number of simultaneously transmitting users that can be served in the same area is also limited. In scenarios with low coherence times, such as the smart traffic scenarios envisioned in 6G, this severely restricts the scalability of the IoT protocols.
The performance of grant-based schemes can be greatly improved by relaxing the orthogonality requirement of the preambles, which significantly increases the number of preambles available within a coherence block. While this can be done without increasing the complexity of the transmitter, it necessitates more sophisticated preamble decoding techniques, such as sparse recovery algorithms that rely on compressed sensing [16, 17]. In this case, the number of available preambles needs to be selected to balance the collision probability, the probability of false alarms, and the complexity of the activity detection algorithm. When the total number of users is not too large, the error is dominated by collisions rather than false alarms, and it may be advantageous to assign a unique preamble to each user to completely avoid collisions. Furthermore, this enables user identification “for free”, since the transmitting users can be identified based from the decoded preambles. By allocating multiple preambles to each device, it is even possible to embed a few bits of information in the preambles as well, e.g., to indicate the amount of data that the user has to transmit [18]. Besides avoiding collisions, the idea of assigning dedicated preambles to each user can be further justified by the result that the probability of activity detection error can theoretically be made to go to zero as the number of base station antennas goes to infinity [17]. However, when the total number of users is large, the computational complexity of the activity detection algorithm may be prohibitive, and false alarms may also be more prominent than collisions. In this case, letting the users randomly select their preamble may lead to better overall performance-complexity tradeoff. We return to the use of sparse recovery for massive uplink in Section II-C.
The main disadvantage of grant-based random access is the latency and power consumption introduced by the three protocol steps. Furthermore, when the packets are short, the additional overhead introduced by the transmission of the grant message may outweigh the advantage of having dedicated channel resources for the packet transmissions. An alternative strategy that avoids this procedure is grant-free random access, which is discussed next.
@@ II-B2 Grant-Free Random Access
In grant-free random access, illustrated in Fig. 3(b), the active devices transmit their preambles and packets immediately, typically in a sequential manner, without waiting for intermediate feedback from the base station. The recently-added Small Data Transmission (SDT) feature in 5G is an example of steps in this direction [19]. To decode the transmitted packets, the base station performs activity detection and channel estimation based on the preambles, and then decodes the messages coherently while treating the interfering transmissions as noise.
@@ II-B3 Unsourced Random Access
The majority of the protocols developed for unsourced random access rely on sparse recovery, possibly combined with additional coding [24, 25], although techniques based on, e.g., tensor decomposition have been proposed as well [26]. Specifically, the work in [24] proposes a scheme where each user transmits its message across multiple sub-blocks. Each sub-block is transmitted non-orthogonally with the other active users, and using sparse recovery techniques the receiver tries to recover the transmissions within each sub-block. Using an outer tree code, the receiver tries to stitch together the decoded transmissions from each sub-block into complete messages. The same idea is adopted in [25], but with an alternative sparse recovery algorithm that is better suited for handling a very large number of preambles. Finally, we note that many random access schemes designed for grant-free access can also be made unsourced by making the preamble selection random or determined based on the message data (as opposed to assigning a unique preamble to each user).
@@ II-C Sparse Recovery for Massive Random Access
As argued in the previous section, many modern random access methods rely on the use of non-orthogonal preambles to enlarge the set of available preambles, so as to reduce the probability of preamble collisions. As a central component towards realizing this goal, this section provides a brief review of the foundations behind sparse recovery for massive random access. To simplify the presentation, we will focus on the case where each user is assigned a unique preamble, which precludes collisions. Clearly, supporting a massive number of users under this strategy is only possible with a very large number of preambles, which necessitates the preambles to be non-orthogonal. We consider a scenario with $N$ single-antenna devices and a single base station with $M$ antennas. Let $\mathbf{s}_{k}\in\mathbb{C}^{L\times 1}$ denote the preamble assigned to user $k$, $k=1,\ldots,N$, which is assumed to be known at the base station. Suppose a random number $K$ out of the $N$ users are active, and let $a_{k}=1$ if user $k$ is active, otherwise $a_{k}=0$. The channel between device $k$ and the base station is assumed to be given by the random vector $g_{k}\mathbf{h}_{k}\in\mathbb{C}^{M\times 1}$, composed by large scale fading coefficient $g_{k}\in\mathbb{R}$, and unknown Rayleigh fading coefficients $\mathbf{h}_{k}\in\mathbb{C}^{M}$ drawn independently from $\mathcal{CN}(0,1)$. The signal $\mathbf{Y}\in\mathbb{C}^{L\times M}$ received by the base station can then be written as